# Sylber: Syllabic Embedding Representation of Speech from Raw Audio

### v1

@@ 5.2 Evaluation
Syllable detection and discovery We evaluate syllable boundaries with precision, recall, F1, and R-value with a 50 ms tolerance, following (Räsänen et al., 2009; Peng et al., 2023; Cho et al., 2024b; Komatsu & Shinozaki, 2024). Syllable discovery is evaluated by a separate clustering, where we use the same process as the previous works that use 4096 clusters. Then, we measure syllable purity, cluster purity, and mutual information between discovered syllables and ground truths (Cho et al., 2024b; Komatsu & Shinozaki, 2024). The same LibriSpeech dev/test sets are used as prior works.
Speech resynthesis We use the extracted speaker embedding and mean pitch from the original speaker to synthesize speech from articulatory features predicted from tokens. We use the original segment durations for each token without predicting them since the duration information can be easily tokenized. (For example, duration can be tagged for each token. See Appendix A.2.) We remove randomness in CFM to yield consistent generation for evaluation purposes. We measure reconstruction performance using the average Pearson Correlation of each component in articulatory features. To evaluate intelligibility, we use an off-the-shelf speech recognition model, Whisper (Radford et al., 2023) [^6] We use “openai/whisper-large-v3” from Huggingface., and measure word error rate (WER) and character error rate (CER). Lastly, we apply an automated speech quality measurement, UTMOS (Saeki et al., 2022), to evaluate the quality of generated speech. These are evaluated on the test-clean split of LibriTTS-R.
@@
Spoken Language Understanding (SLU) To evaluate the language understanding, we use the zero-shot metrics of lexical learning, sWUGGY, and syntax learning, sBLIMP, following Lakhotia et al. (2021); Algayres et al. (2023). These metrics are originally from the Zerospeech Challenge (Nguyen et al., 2020), for discriminating real words/phrases and fake ones. The tasks are solved zero-shot by choosing words/phrases with higher probability inferred by an LM and accuracy is reported.
Categorical perception For the models with frame-wise features, we use dynamic time warping to find an alignment that maximizes similarity. While some sample pairs are already aligned, we find that additional warping yields better scores. For models with syllabic features, we average across all speech (or norm thresholded) parts of the features as all samples are monosyllabic, yielding a single embedding per sample. We use cosine similarity to measure similarity between embeddings from samples, and evaluate DI as defined in §4.
@@ 5.2.1 Baselines
For syllable detection and discovery, we compare our models against HuBERT, VGHuBERT, SDHuBERT, and Komatsu & Shinozaki (2024). For token-to-speech, we train the baseline CFM models using HuBERT units with 50, 100, and 200 cluster sizes as used in Lakhotia et al. (2021), and SDHuBERT tokens with 5K, 10K, and 20K cluster sizes. For coding efficiency, we apply Byte Pair Encoding (BPE) using SentencePiece [^7] https://github.com/google/sentencepiece to merge frequent units to form larger vocabulary that matches ours: 5K, 10K, and 20K, similar to Shen et al. (2024). [^8] The coding efficiency metrics are substantially worse using HuBERT without BPE due to their sampling granularity; thus, we compare against HuBERT with BPE to make a more fair comparison. For evaluating language understanding, we use GSLM (Lakhotia et al., 2021) and tGSLM (Algayres et al., 2023) as baselines. We exclude other uLMs that use more extensive resources than ours (e.g., including a larger dataset or using a pretrained text-based LM.) However, we include tGSLM since it has a similar token granularity (5 Tok/s) as ours although it is trained on a 6$\times$ larger dataset. For phonetic discriminability, we compare Sylber with traditional acoustic features (Melspectrogram and MFCC), representative frame-wise SSL models (HuBERT, Wav2Vec2, and WavLM), and SDHuBERT. For frame-wise SSL models, the best layers with the lowest DIs are chosen.
@@ 6.1 Syllable Detection and Discovery
Table 1: Syllable detection and discovery results. Pr: precision, Re: recall, R: R-value, SP: syllabic purity, CP: cluster purity, and MI: mutual information. Complexity indicates time complexity of post-hoc segmentation algorithm. $n$: the number of frames and $k$: the number of syllables. As we can see, only Sylber uses a linear time algorithm while the other models use a quadratic time algorithm.
@@
Table 1 shows a comparison of syllable detection and discovery performance. Sylber outperforms all previous approaches in every metric other than recall and cluster purity. As these two terms can be inflated by having more segments, it indicates that SDHuBERT is oversegmenting. In terms of discovery, we find the ground truth syllables are more purely mapped to ours than the baselines, greatly improving the previous SOTA by huge margin (59.4 $\rightarrow$ 64.0). The results indicate that our model can detect and discover syllables better than the previous approaches. Moreover, the output features from our model are significantly cleaner than HuBERT or SDHuBERT as shown in Figure 2, showing highly consistent similarities within syllable spans. This allows for a much faster $O(n)$ algorithm applicable, compared to the previous $O(kn^{2})$ and $O(n^{2}/k)$ algorithms where $n$ is the number of frames and $k$ is the estimate number of syllables controlled by a hyperparameter.
@@ 6.2 Resynthesis Performance and Coding Efficiency
The results of token-to-speech resynthesis are shown in Table 2 and can be heard here. We find a general trend in both SDHuBERT and our syllabic tokens that articulatory reconstruction and intelligibility increase with finer clustering granularity. For intelligibility, our model outperforms SDHuBERT at every vocab size while requiring less tokens per second. Interestingly, the articulatory reconstruction is generally higher in SDHuBERT, but also less intelligible. This indicates that our model marginalizes out some amount of articulatory variance which is orthogonal to orthographic contents. This marginalization is also happening in intonation when the embeddings are quantized, as shown in the huge reduction in pitch correlation compared to non-quantized model, resulting in a flattened speech generation. This pattern is shared in both SDHuBERT and our model.
Compared to HuBERT units which have token granularity at sub-phonemic level, the articulation is better reconstructed by HuBERT units with 200 clusters than the units from SDHuBERT or our model. This is natural given their temporal granularity as 28.97 tokens per second, which is likely to capture the local dynamics of articulation better than syllabic level. The intelligibility is also better in the case of 100 and 200 cluster sizes, WERs of 7.78 and 6.34, compared to the best case of syllabic units, WER of 7.95. Though the difference is marginal, HuBERT units require at least 6 times more tokens per second. Also, we find the HuBERT is worse in representing pitch, as suggested by Polyak et al. (2021); Kharitonov et al. (2021). While no significant difference is perceived in quality, our model with 20K vocab size quantization shows the best average UTMOS of 4.21.
@@
To better characterize coding efficiency, we compare bandwidth and coding-rate in Table 3 against baselines with comparable settings. Our model outperforms each baseline in every metric, showing about a 20% gain over the SDHuBERT tokens. In addition, Table 3 demonstrates the innate inefficiency in previous approaches using HuBERT units. There is a minimal gain in sequence compression while increasing the vocabulary size, where BPE is not able to reduce Tok/s by even half of the original when applied to 100 and 200 clusters. The only comparable baseline is BPE on 50 HuBERT clusters, which can reduce Tok/s from 23.59 to between 6.30 and 7.45. However, there is a huge information loss as shown in the high WER of 13.32, which results in a lower coding-rate (0.0283, 0.0285, 0.0287) compared to ours (0.0315, 0.0302, 0.0289) for vocab size of (5K, 10K, 20K) respectively. Without quantization, the best performance is achieved by our model, with a WER of 4.88, Tok/s of 4.27, and higher correlations in loudness and pitch (Table 2). This indicates the significant potential of syllabic tokens as an efficient speech coding that can be harnessed by a better quantization method like vector quantization (VQ) or residual VQ (RVQ), which we leave for future work.
6.3 Spoken Language Understanding (SLU)
Table 4: SLU Results
@@ 6.3 Spoken Language Understanding (SLU)
GSLM 68.70 57.06
tGSLM 68.53 55.31
SDHB-5K 65.80 54.87
SDHB-10K 67.42 54.48
SDHB-20K 67.85 54.87
Sylber-5K 67.32 57.34
Sylber-10K 68.41 58.04
Sylber-20K 70.27 57.67
@@
Table 4 compares the sWUGGY and sBLIMP scores of unit LMs with different vocab size of syllabic units (marked as “model-vocab size”), and the baseline models. Sylber with 20K vocab size outperforms GSLM and tGSLM in sWUGGY, and outperforms baselines at sBLIMP, while none of the SDHuBERT-based LMs outperform GSLM or tGSLM. This indicates that our syllabic tokens have better utility in terms of language modeling compared to the syllabic tokens from SDHuBERT. We also observe a general trend shared in SDHuBERT and ours that a larger vocab size yields a higher sWUGGY score, indicating a finer clustering better covers the lexical space. However, no such pattern was found in sBLIMP score. Notably, Sylber outperforms tGSLM which has a similar token granularity as 5 Hz but with a fixed pooling window, even though their model is trained on a larger dataset. This suggests that a variable pooling window which is dynamically driven by segmentation algorithm is better than using a fixed pooling window.
6.4 Demonstration of Emergent Categorical Perception
Table 5: Phonetic discriminability results. DI: Discriminability Index. Scores for different contrast postitions (onset, coda and all) are separately reported.
@@ 6.4 Demonstration of Emergent Categorical Perception
Frame-wise Wav2Vec2 0.172 0.178 0.174
HuBERT 0.136 0.152 0.141
Wav2Vec2 Large 0.138 0.156 0.143
HuBERT Large 0.166 0.180 0.170
WavLM Large 0.136 0.148 0.140
Syllable-wise SDHuBERT 0.133 0.126 0.131
@@
Figure 3-C illustrates that a clear boundary is drawn when interpolating between two rhyming words, whereas such a boundary is less prominent in HuBERT. This indicates that Sylber needs only 2 categories to represent the interpolating continuum, while HuBERT requires multiple categories or units, which induces high level of inefficiency and redundancy in clustering (Sicherman & Adi, 2023). The curves extracted from the melspectrogram resemble an X-shape, indicating a non-categorical embedding space. This is the closest value to the hypothetical non-categorical score of 0.25 among the models (Table 5). As shown in Table 5, our model’s embeddings demonstrate the best discriminability, with the lowest DIs across both onset and coda contrasts, resulting in an overall DI of 0.112. SDHuBERT also shows lower DIs than frame-wise models, though it still falls short of our model’s performance. The results are highly surprising since we only impose the model to learn temporal structure, and our loss objective does not involve categorical learning at all. However, the embedding space of Sylber is naturally structured to be categorical, indicating the self-segmentation distillation might be a natural learning algorithm that resembles human language learning. Amongst the frame-wise models, WavLM-Large exhibits the best discriminability, which aligns with the fact that WavLM generally has superior representational power compared to other SSL models. Taken together, these qualitative and quantitative results suggest that the embedding space of Sylber is readily quantized, contributing to the performance improvements observed in previous sections.

### v2

@@ 4.2 Evaluation
Syllable detection and discovery We evaluate syllable boundaries with precision, recall, F1, and R-value following previous studies (Peng et al., 2023; Cho et al., 2024b; Komatsu & Shinozaki, 2024). Syllable discovery is evaluated by a separate clustering, where we use the same process as the previous works that use 4096 clusters. We measure syllable purity, cluster purity, and mutual information between the discovered syllables and the ground truth (Cho et al., 2024b; Komatsu & Shinozaki, 2024). See Appendix A.1.8 for details.
Speech resynthesis We measure reconstruction performance using the average Pearson Correlation of each component in articulatory features. To evaluate intelligibility, we use an off-the-shelf speech recognition model, Whisper (Radford et al., 2023) [^4] We use “openai/whisper-large-v3” from Huggingface., and measure word error rate (WER) and character error rate (CER). Lastly, we apply an automated speech quality measurement, UTMOS (Saeki et al., 2022), to evaluate the quality of generated speech. These are evaluated on the test-clean split of LibriTTS-R. For some models, we collect subjective human evaluation on qualities, and report mean opinion scores (MOS) on the naturalness (nMOS) and prosodic similarity with the ground truth (psMOS).
@@
Spoken Language Understanding We use the zero-shot metrics of lexical learning, sWUGGY, and syntax learning, sBLIMP, following Lakhotia et al. (2021); Algayres et al. (2023). These metrics are measured by accuracy of discriminating real words/phrases and fake ones using the probabilities inferred from uLM, respectively.
@@ 4.2.1 Baselines
For syllable detection and discovery, we compare our models against HuBERT, VGHuBERT, SDHuBERT, and Komatsu & Shinozaki (2024). For token-to-speech, we train the baseline CFM models using deduplicated HuBERT units with the size of 50, 100, and 200 by Lakhotia et al. (2021), and 500, and 2K by Nguyen et al. (2023); and SDHuBERT tokens with 5K, 10K, and 20K cluster sizes. For coding efficiency, we apply Byte Pair Encoding (BPE) using SentencePiece [^5] https://github.com/google/sentencepiece to merge frequent units to form a larger vocabulary for HuBERT so that the size matches ours: 5K, 10K, and 20K, similar to Shen et al. (2024). [^6] The coding efficiency metrics are substantially worse using HuBERT without BPE due to their sampling granularity; thus, we compare against HuBERT with BPE to make a more fair comparison. For evaluating language understanding, we use GSLM (Lakhotia et al., 2021), tGSLM (Algayres et al., 2023), NAST (Messica & Adi, 2024), TWIST (Hassid et al., 2024) as baselines. These are selected as their tokenizers stem from HuBERT.
@@ 5.1 Syllable Detection and Discovery
Table 1: Syllable detection and discovery results measured. Pr: precision, Re: recall, R: R-value, SP: syllabic purity, CP: cluster purity, and MI: mutual information. Complexity indicates time complexity of post-hoc segmentation algorithm. $n$: the number of frames and $k$: the number of syllables. All metrics are measured using LibriSpeech test data. Results of other model–algorithm combinations are also compared. Sylber uses a linear time algorithm while the other models use a quadratic time algorithm, which is only available with the clean structure learned by our approach.
@@
Sylber–MinCut $O(n^{2}/k)$ 76.8 68.1 72.2 75.8 63.9 44.0 5.29
HuBERT-Greedy $O(n)$ 54.5 35.2 42.8 52.7 29.5 25.9 3.36
SDHuBERT-Greedy $O(n)$ 56.1 67.4 61.2 62.1 30.0 41.5 2.67
Table 2: Syllable detection performance in out-of-domain data. Sylber can generalize across novel domain and langauges without any tuning, showing high scores in all detection metrics.
[…]
LibriSpeech (in-domain) English Reading 76.6 68.3 72.2 75.9
Fisher English Conversation 78.8 66.2 71.9 75.0
MLS Spanish Reading 73.5 69.9 71.7 75.9
AISHELL-3 Mandarin Reading 74.9 68.0 71.3 75.3
Table 1 shows a comparison of syllable detection and discovery performance. Sylber outperforms all previous approaches in every metric other than recall and cluster purity. As these two terms can be inflated by having more segments, it indicates that SDHuBERT is oversegmenting. In terms of discovery, we find the ground truth syllables are more purely mapped to ours than the baselines, greatly improving the previous SOTA by a huge margin (59.4 $\rightarrow$ 64.0). The results indicate that our model can detect and discover syllables better than the previous approaches. (Check Appendix A.2.7 for visualizations of the embedding space.) Moreover, the output features from our model are significantly cleaner than HuBERT or SDHuBERT as shown in Figure 2, which shows highly consistent similarities within syllable spans. This allows for a much faster $O(n)$ algorithm to be applicable, compared to the previous $O(kn^{2})$ and $O(n^{2}/k)$ algorithms where $n$ is the number of frames and $k$ is the estimated number of syllables controlled by a hyperparameter. Compared to SDHuBERT, our syllable segmentation shows approximately 4$\times$ gains in inference time (Appendix A.2.8).
When we apply the previous algorithm, MinCut (Peng et al., 2023; Cho et al., 2024b) to Sylber, the scores show very marginal differences (Sylber-Mincut in Table 1). MinCut is designed to search for optimal segments at the cost of computation time. Therefore, this indicates that Sylber features are clean and robust enough to find optimal segments in a greedy manner. In fact, when the greedy linear-time algorithm is applied to SDHuBERT, we find significant performance degradation (SDHuBERT-Greedy). In the case of HuBERT, both Greedy and MinCut show generally low performance due to the lack of syllabic structure.
Moreover, Sylber demonstrates surprising generalizability to unseen domain and languages without any tuning. As shown in Table 2, the syllable detection by Sylber shows high performance when applied to other corpora: conversational speech (Fisher; Cieri et al. (2004)), Spanish (MLS; Pratap et al. (2020)), and Mandarin (AISHELL-3; Shi et al. (2021)), which are on par with the in-domain case (LibriSpeech). This multilingual generalizability indicates that Sylber represents phonological information which may have a shared physical basis across languages (Ohala, 1984; 1990; Cho et al., 2024a). This finding suggests a potential scalability of Sylber. More details and discussion are in Appendix A.2.5.
@@ 5.2 Resynthesis Performance and Coding Efficiency
The results of token-to-speech resynthesis are shown in Table 3 and can be heard here. We find a general trend in both SDHuBERT and our syllabic tokens that articulatory reconstruction and intelligibility increase with finer clustering granularity. For intelligibility, our model outperforms SDHuBERT at every vocab size while requiring less tokens per second. Interestingly, the articulatory reconstruction is generally higher in SDHuBERT, but also less intelligible. This indicates that our model marginalizes out some amount of articulatory variance which is orthogonal to orthographic contents. This marginalization is also happening in intonation when the embeddings are quantized, as shown in the huge reduction in pitch correlation compared to non-quantized model, resulting in a flattened speech generation. This pattern is shared in both SDHuBERT and Sylber.
The articulation is slightly better reconstructed by HuBERT units with 100 or more clusters than the units from SDHuBERT or our model, as shown in Table 3. This can be attributed to HuBERT’s sub-phonemic, articulatory representations (Cho et al., 2023), which have a fine temporal granularity of 26.68 Tok/s or higher. The intelligibility is also slightly better with HuBERT units of 100 or more clusters, with WERs of 5.04–7.78, compared to the best-performing syllabic units with a WER of 7.95. Though the difference is marginal, HuBERT units require at least 6-8 times more tokens per second. Also, we find HuBERT units perform worse in representing pitch, corroborating findings by Polyak et al. (2021); Kharitonov et al. (2021); Nguyen et al. (2023).
To further evaluate coding efficiency, we compare Sylber against baselines with comparable settings of HuBERT units in Table 4. Our model outperforms each baseline in every metric, showing about a 20% gain over the SDHuBERT tokens. In addition, Table 4 demonstrates the innate inefficiency in previous approaches using HuBERT units. There is a minimal gain in sequence compression while increasing the vocabulary size, where BPE is not able to reduce Tok/s by even half of the original when applied to 100 and 200 clusters. The only comparable baseline is BPE on 50 HuBERT clusters, which can reduce Tok/s from 23.59 to between 6.30-7.45. However, there is a huge information loss as shown in the high WER of 13.32, which results in a lower coding-rate (0.0283, 0.0285, 0.0287) compared to ours (0.0315, 0.0302, 0.0289) for vocab size of (5K, 10K, 20K) respectively.
Table 5: Subjective evaluation on resynthesis quality.
@@
GT 4.37 4.71
HB 200 3.24 2.65
HB 2K 3.33 2.90
Sylber 20K 3.32 3.04
Sylber $\infty$ 3.80 3.62
HuBERT units and Sylber units show comparable quality in terms of naturalness in both machine and human evaluation (UTMOS in Table 3; nMOS in Table 5). In terms of prosody, Sylber 20K units show higher subjective similarity than HuBERT 200 or 2K units as shown in the psMOS results (Table 5).
Without quantization, the best performance is achieved by our model, with a WER of 4.88, Tok/s of 4.27, and higher correlations in loudness and pitch (Table 3). Furthermore, both of the subjective qualities, nMOS and psMOS, significantly increase (Table 5). This indicates the significant potential of syllabic tokens as an efficient speech coding that can be harnessed by a better quantization method. We leave this investigation for future work.
5.3 Spoken Language Understanding
Table 6: Speech uLM performance comparison. Sections are divided by training data size (top: LibriSpeech (LS) and bottom: LibriLight (LL) or more).
[…]
GSLM 150M 100 26.68 177.26 LS 1K 68.70 57.06
SDHuBERT-uLM 125M 5K 5.24 64.39 LS 1K 65.80 54.87
10K 69.63 LS 1K 67.42 54.48
20K 74.87 LS 1K 67.85 54.87
Sylber-uLM 125M 5K 4.27 52.47 LS 1K 67.32 57.34
10K 56.74 LS 1K 68.41 58.04
20K 61.01 LS 1K 70.27 57.67
tGSLM 150M 5 LL 6K 68.53 55.31
NAST 150M 200 28.97 221.44 LL 6K 76.42 55.62
TWIST-ColdInit 125M 500 16.78 150.45 LL++ 150K 77.74 54.27
TWIST 13B 500 16.78 150.45 LL++ 150K 84.10 59.20
Sylber-uLM 125M 20K 4.27 61.01 LL 66K 76.31 60.54
Sylber-w/SIL-uLM 125M 20K 4.76 68.01 LL 66K 78.03 60.78
@@ 5.3 Spoken Language Understanding
Table 6 compares the sWUGGY and sBLIMP scores of speech uLMs. In the limited resource setting that uses 1K hours of training data. Sylber-uLMs generally outperform the baselines, GSLM and SDHuBERT-uLMs at sBLIMP, and the model with 20K vocab size outperforms GSLM in sWUGGY, while none of the SDHuBERT-uLMs outperform GSLM or Sylber-uLMs (top section of Table 6). This indicates that our syllabic tokens have better utility in terms of language modeling compared to the syllabic tokens from SDHuBERT. We also observe a general trend shared in SDHuBERT and ours that a larger vocab size yields a higher sWUGGY score, indicating that a finer clustering better covers the lexical space. Notably, Sylber-uLM trained on 1K hours outperforms tGSLM which has a similar token granularity as 5 Hz but with a fixed pooling window, even though their model is trained on a larger dataset of 150K hours. This suggests that a variable pooling window by syllabic segmentation is more suitable than using a fixed pooling window.
When we scale up the train data to 66K hours, Sylber-uLMs are able to achieve comparable or better sWUGGY scores than the models with similar sizes: tGSLM, NAST, and TWIST-ColdInit (bottom section of Table 6). We also include a silence interleaved version (Sylber-w/SIL-uLM), where we insert a silence token when the gap between two adjacent tokens is longer than 140 ms. This improves the performance, especially sWUGGY, while increasing Tok/s and bitrate to 4.76 and 68.01, which are still far below the previous models. Although sWUGGY falls short of the TWIST model with 13B parameters, Sylber-uLMs show slightly better performance in sBLIMP, which is astonishing given the huge gap in model size and training data. Most importantly, all these results are obtained using significantly smaller sequence length and bitrate, several times lower than previous approaches. This suggests that Sylber units are highly efficient and valid tokens for spoken language modeling.
6 Emergent Categorical Perception in Sylber
Figure 3: A. Overview of articulatory interpolation of rhyming words when interpolating $\alpha\in[0,1]$. B. Hypothetical curves of categorical (solid lines) and non-categorical (dashed lines) embeddings. C. Similarity curves examples from Melspectrogram (Mel), HuBERT, and Sylber. Sylber consistently shows highly categorical perception, drawing a sharp boundary in continuum between words.
The results shown in Section 5.2 suggest that HuBERT features might densely tile the embedding space, requiring many clusters to fully cover the space with fidelity. This lack of discreteness leads to a redundant token vocabulary (Sicherman & Adi, 2023), resulting in an exponentially growing token vocabulary size when tokens are grouped to reduce sequence length. However, human speech perception exhibits categorical effects, where listeners tend to impose discrete boundaries on a continuum of speech sounds rather than tracking gradual acoustic changes (Liberman et al., 1957; Pisoni & Lazarus, 1974; Harnad, 2003). This phenomenon—categorical perception—suggests that phonemic categories are perceived as discrete units despite underlying acoustic variability. If similar categorical effects emerge in the embedding space of SSL models, they could enable more efficient tokenization with a compressed set of discrete units. Based on this linguistic theory, we evaluate the degree of categorical perception in Sylber and other SSL models to assess the discreteness of their embedding spaces.
We simulate interpolation between two rhyming words to probe the embeddings of speech SSL models to check whether they are categorical. Specifically, monosyllabic words are used where a single consonant is different at the front (onset) or back (coda) of the syllable. We make the contrast to be varying one of the phonological properties: nasality, voicedness, or place (e.g., “ball” vs “mall”, “down” vs “town”, or “lest” vs “rest”, respectively). We do not include vowel contrasts since categorical perception of vowels is not as consistent as consonants (Pisoni, 1973; Pisoni & Lazarus, 1974). We consider 13 types of such difference and simulate 4 pairs for each type, resulting 52 word pairs in total. The details of the difference types and the full list of word pairs can be found in Appendix A.1.9. To simulate a continuum between words, we utilize SPARC (Cho et al., 2024c) which allows direct editing in the physical articulatory space (Figure 3-A). We first generate audio using the an off-the-shelf TTS API. [^7] We use the TTS service in Vertex AI (https://cloud.google.com/vertex-ai) with a default female voice. We extract articulatory features from the speech, which are then temporally aligned by dynamic time warping to either end if necessary. We sample 51 equidistant samples in the linear interpolation between two words, where each end is manually adjusted to make the perceptual boundary drawn approximately in the middle ($\alpha=0.5$), which can be heard here. The pitch and loudness are also controlled to be at the same level. More details about SPARC can be found in Appendix A.1.4.
Given a speech representation model, we extract features for each interpolating point between words in each pair. We calculate the similarity between interpolating features with features from either end, forming a likelihood curve along the interpolation. Hypothetically, if the representation is categorical, the likelihood curves should show a sharp transition at the boundary (Figure 3-B). If the embeddings are not categorical and tracing the interpolation, the curves would show “X” pattern as the dashed line in Figure 3-B. We define the Discriminability index (DI) to quantify the level of categorical perception. The DI measures an empirical risk of wrong discrimination, where the probability is calculated based on similarities. See Appendix A.1.10 for the detailed definition. If the embeddings are categorical, DI will be close to 0. If they are non-categorical with X-shaped curves, DI will be 0.25. The maximum value of DI is 0.5, which would be random chance discrimination.
Figure 3-C illustrates that a clear boundary is drawn when interpolating between two rhyming words, whereas such a boundary is less prominent in HuBERT. This indicates that Sylber needs only 2 categories to represent the interpolating continuum, while HuBERT requires multiple categories or units, which induces high levels of inefficiency and redundancy. The similarity curves using the melspectrogram resemble an X-shape, indicating a non-categorical embedding space.
Table 7: DI comparison.
@@ 6 Emergent Categorical Perception in Sylber
W2V2 0.172 0.178 0.174
HB 0.136 0.152 0.141
W2V2-L 0.138 0.156 0.143
HB-L 0.166 0.180 0.170
WavLM-L 0.136 0.148 0.140
SDHB 0.133 0.126 0.131
@@
We compare Sylber with traditional acoustic features (Melspectrogram and MFCC), representative frame-wise SSL models (HuBERT (HB), Wav2Vec2 (W2V2), WavLM, and large (-L) versions (if applicable), and SDHuBERT. As shown in Table 7, our model’s embeddings demonstrate the best discriminability, with the lowest DIs across both onset and coda contrasts (overall DI: 0.112). The results are unexpected as we only impose the model to learn temporal structure, and our loss objective does not involve any categorical learning at all. However, the embedding space of Sylber is naturally structured to be categorical, indicating the self-segmentation distillation might be a natural learning algorithm that resembles human language learning. Taken together, these qualitative and quantitative results suggest that the embedding space of Sylber is readily quantized, contributing to the performance improvements observed in previous sections.